• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin Huang

    Log entry time 08:24:38 on February 15, 2009

    Entry number 261267

    This entry is a followup to: 261265

    keyword=Access Summary: Problem with BB Trigger

    Brad and I took a control access from 1:16AM to 5:34AM to inspect trigger non-uniformity on BB spectrometer. Several problems with BigBite trigger have been identified, which could impact both d2n and transversity.


    Following is a preliminary description; Brad is going to bring more details.



    1st Problem:

    One of the NIM bins, fully installed with BB shower summing modules, is overloaded on its +-12V power supply. This would reduce the shower signal gain by ~1 order of magnitude to both Trigger and DAQ, as well as introduce distortion.



    If I understand right, this NIM bin is holding summing modules, except shower row 24 25,26,27(which are in another normal working bin). When Brad and I first looked at this bin, its -12V is only outputting -1.5V. During follow tests, its +12V power also dropped to +1.5V in some case.


    For our test signal sent to one of the input:


    Power supply State

    Back Panel

    output (To ADC)

    Front Panel Sum Output

    (To Triggers)


    Normal Voltages



    Signal Shape is Normal

    -12V Failure



    Signal is also distorted

    +12V Failure



    Signal Shape is Normal


    It’s also observed that, if a test signal is sent at PMT signal cable, the read out at T1 trigger discriminator inputs are:

    ·         -70-90mV for sum#1->15; 17->23; signal is distorted

    ·         0mV for sum#16, see 2nd problem

    ·         -1080mV for sum#24- >26

    This could explain “pion punch through at higher BB acceptance” during Transversity. That could be caused by that shower block row 1->23 is not properly amplified (and for broken channels, its ADC signal copy could be amplified more than its trigger discriminator copy).


    Temp. Solution:

    4 of the summing modules are moved to a new NIM crate, which bring power load to normal. Now gains of row 1-23 summing module are back to normal.

    Shower HV gain might need to be recalibrated; threshold need to be reset. We also need to assess its impact to transversity data.



    2nd Problem:

    Fan output problem with shower sum #16.

    1.     The fan base line is not zero-able (we tried to lower it from +175mV -> +75mV, cannot lower more).

    2.     When we try to send a pulse from PMT signal cable, the discriminator of T1 shower sum #16 do not get any signal (other channels are OK except 1st problem)! This could form a trigger whole for Transversity Coin set up (to be confirmed)

    3.     Another fan copy to T6 is OK.

    Since T1 is not used, we left it as it was.



    3rd Problem:

    T1 trigger, Shower sum #19 has a flippy cable. Replaced by Brad



    4th  Problem:

    Brad: “P-26L” 50 Ohm splitter is unplugged”; should be fixed by him.

    A copy of this log entry has been emailed to: brads, kalyan, moffit, xqian, jinhuang